# A Class of Generalized Shift-Splitting Preconditioners for Double Saddle Point Problems

### v3

@@ 3 Proposed generalized shift-splitting (GSS) iterative method and preconditioner
This section proposes a generalized shift-splitting (GSS) iterative method to solve the DSPP. Let $\alpha,\beta,\tau,$ $\bm{\omega}$ be positive real numbers and $P\in{\mathbb{R}}^{n\times n},Q\in{\mathbb{R}}^{l\times l},$ and $R\in{\mathbb{R}}^{m\times m}$ be SPD matrices, then the saddle point $\mathcal{B}$ admits the following matrix splitting:
$\mathcal{B}=(\Theta+{\bm{\omega}}\mathcal{B})-(\Theta-(1-\bm{\omega})\mathcal{B})=:\mathcal{M}_{\mathrm{GSS}}-\mathcal{N}_{\mathrm{GSS}},$ (3.1)
@@
$\mathcal{M}_{\mathrm{GSS}}=\left[\begin{matrix}\alpha P+\bm{\omega}A&\bm{0}&\bm{\omega}B^{T}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{\omega}C\\ -\bm{\omega}B&-\bm{\omega}C^{T}&\tau R\end{matrix}\right],$ (3.2)
@@
and $\Theta=\left[\begin{matrix}\alpha P&\bm{0}&\bm{0}\\ \bm{0}&\beta Q&\bm{0}\\ \bm{0}&\bm{0}&\tau R\end{matrix}\right].$
@@
where ${\bm{u}}^{(k)}=\left[{x^{(k)}}^{T},{y^{(k)}}^{T},{z^{(k)}}^{T}\right]^{T}\in{\mathbb{R}}^{n+l+m},$ $\mathcal{G}=\mathcal{M}^{-1}_{\mathrm{GSS}}\mathcal{N}_{\mathrm{GSS}}$ is the iteration matrix and ${\bm{d}}=\mathcal{M}_{\mathrm{GSS}}^{-1}{\bm{b}}\in{\mathbb{R}}^{n+m+l}.$
The matrix splitting in (3.1) induces a preconditioner, denoted as $\mathcal{M}_{\mathrm{GSS}}$, which can be utilized to speed up the convergence rate of the Krylov subspace iterative methods, like GMRES. This preconditioner is referred to as the GSS preconditioner.
@@
$\mathcal{M}_{\mathrm{\mathrm{GSS}}}{\bm{z}}=r,$ (3.5)
@@
$\mathcal{M}_{\mathrm{GSS}}=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -\bm{\omega}B(\alpha P+\bm{\omega}A)^{-1}&-\bm{\omega}C^{T}(\beta Q+\bm{\omega}D)^{-1}&I\end{matrix}\right]\left[\begin{matrix}\alpha P+\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\widehat{R}\end{matrix}\right]$
.$\hskip 113.81102pt.\left[\begin{matrix}I&\bm{0}&\bm{\omega}(\alpha P+\bm{\omega}A)^{-1}B^{T}\\ \bm{0}&I&\bm{\omega}(\beta Q+\bm{\omega}D)^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$
@@
Algorithm 1 Solving $\mathcal{M}_{\mathrm{GSS}}{\bm{z}}={\bm{r}}$
Input: The matrices $A,B,C,D,$ $\bm{r},$ positive parameters $\alpha,\beta,\tau,\bm{\omega},$ SPD matrices $P\in{\mathbb{R}}^{m\times m},$ $Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m}.$
@@
Algorithm 1 necessitates solving two linear subsystems having the coefficient matrix $(\alpha P+\omega A)$, two subsystems having the coefficient matrix $(\beta Q+\bm{\omega}D),$ and one subsystem with the coefficient matrix $\widehat{R}.$ Since the coefficient matrices are SPD matrices, we have the flexibility to employ exact solvers, such as Cholesky factorization, and inexact solvers, like the preconditioned conjugate gradient method, to solve them efficiently. Nevertheless, to solve steps $1$ and $4$, only one Cholesky factorization of $\alpha P+\bm{\omega}A$ and to solve steps $2$ and $5$, only one Cholesky factorization of $\beta Q+\bm{\omega}D$ is needed to perform.
@@ 4 Convergence analysis of the GSS iterative method
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ are SPD matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{l\times m}$ are full row matrices. Then the matrix $\Theta^{-1}\mathcal{B}$ is positive stable.
@@
Assume that $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ are SPD matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{l\times m}$ are full row matrices. Let $\alpha,\beta,\tau,\bm{\omega}>0,$ $P\in{\mathbb{R}}^{n\times n},Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m}$ be SPD matrices. Then, the GSS iterative method converges to the unique solution of the DSPP (1.1) if
@@
$\mathcal{G}=\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{N}_{\mathrm{GSS}}$ $=(\Theta+\bm{\omega}\mathcal{B})^{-1}(\Theta-(1-\bm{\omega})\mathcal{B})$ (4.1)
$=(I-\bm{\omega}\Theta^{-1}\mathcal{B})^{-1}(I-(1-\bm{\omega})\Theta^{-1}\mathcal{B}).$ (4.2)
@@
$(I-\bm{\omega}\Theta^{-1}\mathcal{B})^{-1}\left(I-(1-\bm{\omega})\Theta^{-1}\mathcal{B}\right){\bf x}=\lambda{\bf x},$ (4.3)
@@
$(I-(1-\bm{\omega})\Theta^{-1}\mathcal{B}){\bf x}=\lambda(I-\bm{\omega}\Theta^{-1}\mathcal{B}){\bf x}$ (4.4)
@@
Notably, solving the DSPP (1.1) is same as to solve the preconditioned linear system $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B}{\bf u}=\mathcal{M}_{\text{GSS}}^{-1}{\bm{b}}$. Hence, as an immediate consequence of Theorem 4.1 and Remark 4.1, we have the following results regarding the clustering properties of the spectrum of the preconditioned matrix $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B}$.
@@
Assume that $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{p\times p}$ are SPD matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are full row matrices. Let $\mathcal{M}_{\mathrm{GSS}}$ be defined as in (3.2) and $\lambda$ be an eigenvalue of the preconditioned matrix $\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{B}$. If $\bm{\omega}\geq 1/2,$ then $\lambda$ satisfies the following:
@@
i.e., all eigenvalues of the preconditioned matrix $\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{B}$ are entirely contained in a circle centered at $(1,0)$ with radius strictly less one.
@@
The proof follows immediately from the identity $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B}=\mathcal{M}_{\text{GSS}}^{-1}(\mathcal{M}_{\text{GSS}}-\mathcal{N}_{\text{GSS}})=I-\mathcal{G}.$blacksquare
@@ 5 Two relaxed variants of GSS preconditioner
To enhance efficiency, this section introduces two relaxed variants of the GSS preconditioner. By removing $\alpha P$ from $(1,1)$ block and $\beta Q$ from $(2,2)$ block of $\mathcal{M}_{\mathrm{GSS}},$ we obtain the following two relaxed GSS (RGSS) preconditioners:
$\mathcal{M}_{\text{RGSS-I}}=\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{\omega}B^{T}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{\omega}C\\ -\bm{\omega}B&-\bm{\omega}C^{T}&\tau R\end{matrix}\right]$ (5.1)
@@
$\mathcal{M}_{\text{RGSS-II}}=\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{\omega}B^{T}\\ \bm{0}&\bm{\omega}D&\bm{\omega}C\\ -\bm{\omega}B&-\bm{\omega}C^{T}&\tau R\end{matrix}\right].$ (5.2)
In the implementation, at each step of RGSS preconditioners in conjunction with GMRES, we are required to solve the linear systems of the following forms:
$\mathcal{M}_{\text{RGSS-I}}{\bm{w}}={\bm{r}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{M}_{\text{RGSS-II}}{\bm{w}}={\bm{r}}.$ (5.3)
@@
$\mathcal{M}_{\text{RGSS-I}}=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -BA^{-1}&-\bm{\omega}C^{T}(\beta Q+\bm{\omega}D)^{-1}&I\end{matrix}\right]\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\mathcal{R}_{1}\end{matrix}\right]\left[\begin{matrix}I&\bm{0}&A^{-1}B^{T}\\ \bm{0}&I&\bm{\omega}(\beta Q+\bm{\omega}D)^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$ (5.4)
$\mathcal{M}_{\text{RGSS-II}}=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -BA^{-1}&-C^{T}D^{-1}&I\end{matrix}\right]\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\mathcal{R}_{2}\end{matrix}\right]\left[\begin{matrix}I&\bm{0}&A^{-1}B^{T}\\ \bm{0}&I&D^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$ (5.5)
By applying a similar methodology as in Algorithm 1, we derive Algorithms 2 and 3 for implementing the RGSS preconditioners.
Algorithm 2 Solving $\mathcal{M}_{\text{RGSS-I}}{\bm{w}}={\bm{r}}$
Input: The matrices $A,B,C,D,$ $\bm{r},$ positive parameters $\beta,\tau,\bm{\omega},$ SPD matrices $Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m}.$
@@
$1:$ Solve $At_{1}=r_{1}/\bm{\omega}$ to find $t_{1}.$
$2:$ Solve $(\beta Q+\bm{\omega}D)t_{2}=r_{2}$ to find $t_{2}.$
$3:$ Solve $\mathcal{R}_{1}\bm{w}_{3}=r_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2}$ to obtain $\bm{w}_{3}.$
$4:$ Solve $A\bm{w}_{1}=\frac{1}{\bm{\omega}}(r_{1}-B^{T}\bm{w}_{3})$ to find $\bm{w}_{1}.$
$5:$ Solve $(\beta Q+\bm{\omega}D)\bm{w}_{2}=r_{2}-\bm{\omega}C\bm{w}_{3}$ to obtain $\bm{w}_{2}.$
Algorithm 3 Solving $\mathcal{M}_{\text{RGSS-II}}{\bm{w}}={\bm{r}}$
Input: The matrices $A,B,C,D,$ $\bm{r},$ positive parameters $\tau,\bm{\omega}$ and the SPD matrix $R\in{\mathbb{R}}^{m\times m}.$
@@
$1:$ Solve $At_{1}=r_{1}/\bm{\omega}$ to find $t_{1}.$
$2:$ Solve $Dt_{2}=r_{2}/\bm{\omega}$ to find $t_{2}.$
$3:$ Solve $\mathcal{R}_{2}\bm{w}_{3}=r_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2}$ to obtain $\bm{w}_{3}.$
$4:$ Solve $A\bm{w}_{1}=\frac{1}{\bm{\omega}}(r_{1}-B^{T}\bm{w}_{3})$ to find $\bm{w}_{1}.$
$5:$ Solve $D\bm{w}_{2}=\frac{1}{\bm{\omega}}(r_{2}-\bm{\omega}C\bm{w}_{3})$ to obtain $\bm{w}_{2}.$
Next, we investigate the spectral properties of the preconditioned matrices $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}.$

### v4

@@ 3 Proposed generalized shift-splitting (GSS) iterative method and preconditioner
This section proposes a generalized shift-splitting (GSS) iterative method to solve the DSPP. Let $\alpha,\beta,\tau,$ $\bm{\omega}$ be positive real numbers and $P\in{\mathbb{R}}^{n\times n},Q\in{\mathbb{R}}^{l\times l},$ and $R\in{\mathbb{R}}^{m\times m}$ be SPD matrices, then the saddle point matrix $\mathcal{B}$ admits the following matrix splitting:
$\mathcal{B}=(\Theta+{\bm{\omega}}\mathcal{B})-(\Theta-(1-\bm{\omega})\mathcal{B})=:{\mathcal{P}}_{\mathrm{GSS}}-\mathcal{N}_{\mathrm{GSS}},$ (3.1)
@@
${\mathcal{P}}_{\mathrm{GSS}}=\left[\begin{matrix}\alpha P+\bm{\omega}A&\bm{0}&\bm{\omega}B^{T}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{\omega}C\\ -\bm{\omega}B&-\bm{\omega}C^{T}&\tau R\end{matrix}\right],$ (3.2)
@@
and
$\Theta=\left[\begin{matrix}\alpha P&\bm{0}&\bm{0}\\ \bm{0}&\beta Q&\bm{0}\\ \bm{0}&\bm{0}&\tau R\end{matrix}\right].$
@@
where ${\bm{u}}^{(k)}=\left[{x^{(k)}}^{T},{y^{(k)}}^{T},{z^{(k)}}^{T}\right]^{T}\in{\mathbb{R}}^{n+l+m},$ $\mathcal{G}={\mathcal{P}}^{-1}_{\mathrm{GSS}}\mathcal{N}_{\mathrm{GSS}}$ is the iteration matrix and ${\bm{d}}={\mathcal{P}}_{\mathrm{GSS}}^{-1}{\bm{b}}\in{\mathbb{R}}^{n+m+l}.$
The matrix splitting in (3.1) induces a preconditioner, denoted as ${\mathcal{P}}_{\mathrm{GSS}}$, which can be utilized to speed up the convergence rate of the Krylov subspace iterative methods, like GMRES. This preconditioner is referred to as the GSS preconditioner.
@@
${\mathcal{P}}_{\mathrm{\mathrm{GSS}}}{\bm{z}}=r,$ (3.5)
@@
${\mathcal{P}}_{\mathrm{GSS}}=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -\bm{\omega}B(\alpha P+\bm{\omega}A)^{-1}&-\bm{\omega}C^{T}(\beta Q+\bm{\omega}D)^{-1}&I\end{matrix}\right]\left[\begin{matrix}\alpha P+\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\widehat{R}\end{matrix}\right]$
.$\hskip 113.81102pt.\left[\begin{matrix}I&\bm{0}&\bm{\omega}(\alpha P+\bm{\omega}A)^{-1}B^{T}\\ \bm{0}&I&\bm{\omega}(\beta Q+\bm{\omega}D)^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$ (3.6)
@@
Algorithm 1 Solving ${\mathcal{P}}_{\mathrm{GSS}}{\bm{z}}={\bm{r}}$
Input: The matrices $A\in{\mathbb{R}}^{n\times n},\,B\in{\mathbb{R}}^{m\times n},\,C\in{\mathbb{R}}^{l\times m},\,D\in{\mathbb{R}}^{l\times l},$ $\bm{r}\in{\mathbb{R}}^{n+l+m},$ positive parameters $\alpha,\beta,\tau,\bm{\omega},$ SPD matrices $P\in{\mathbb{R}}^{m\times m},$ $Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m}.$
@@
Algorithm 1 necessitates solving two linear subsystems having the coefficient matrix $(\alpha P+\omega A)$, two subsystems having the coefficient matrix $(\beta Q+\bm{\omega}D),$ and one subsystem with the coefficient matrix $\widehat{R}.$ We can use LU factorization to solve them efficiently. Moreover, when $A$ and $D$ are SPD matrices, we have the flexibility to employ exact solvers, such as Cholesky factorization, and inexact solvers, like the preconditioned conjugate gradient method, to solve them efficiently. Nevertheless, to solve steps $1$ and $4$, only one Cholesky or LU factorization of $\alpha P+\bm{\omega}A$ and to solve steps $2$ and $5$, only one Cholesky or LU factorization of $\beta Q+\bm{\omega}D$ is needed to perform.
Remark 3.2.
From Algorithm 1, observe that the most tedious task to implement the GSS preconditioner is to solve the linear subsystem in step $3$ To avoid the direct construction of the matrices $B(\alpha P+\bm{\omega}A)^{-1}B^{T}$ and $C^{T}(\beta Q+\bm{\omega}D)^{-1}C,$ we modify the decomposition in (3) in the following way:
${\widetilde{\mathcal{P}}}_{\mathrm{GSS}}:=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -\bm{\omega}B(\alpha P+\bm{\omega}A)^{-1}&-\bm{\omega}C^{T}(\beta Q+\bm{\omega}D)^{-1}&I\end{matrix}\right]\left[\begin{matrix}\alpha P+\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\widetilde{P}+\widetilde{Q}\end{matrix}\right]$
.$\hskip 113.81102pt.\left[\begin{matrix}I&\bm{0}&\bm{\omega}(\alpha P+\bm{\omega}A)^{-1}B^{T}\\ \bm{0}&I&\bm{\omega}(\beta Q+\bm{\omega}D)^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$ (3.7)
where $\widetilde{P}$ and $\widetilde{Q}$ are (efficient and economical) approximations of the matrices $B(\alpha P+\bm{\omega}A)^{-1}B^{T}$ and $C^{T}(\beta Q+\bm{\omega}D)^{-1}C,$ respectively. Then, the step $4$ in Algorithm 1 changes to the linear subsystem $(\widetilde{P}+\widetilde{Q})\bm{z}_{3}={\bm{r}}_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2}.$ With suitably chosen $\widetilde{P}$ and $\widetilde{Q}$ (see for Example $1$), this subsystem is much easier to implement than step $3$ of Algorithm 1. We denote this inexact version of the GSS preconditioner as $\widetilde{P}_{\text{GSS}}.$
@@ 4 Convergence analysis of the GSS iterative method
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ be such that $A_{H}$ and $D_{H}$ positive definite matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{l\times m}$ be full row matrices. Then the matrix $\Theta^{-1}\mathcal{B}$ is positive stable.
@@
Assume that $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ are nonsingular matrices with $A_{H}$ and $D_{H}$ positive definite, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{l\times m}$ are full row matrices. Let $\alpha,\beta,\tau,\bm{\omega}>0,$ $P\in{\mathbb{R}}^{n\times n},Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m}$ be SPD matrices. Then, the GSS iterative method converges to the unique solution of the DSPP (1.1) if
@@
$\mathcal{G}={\mathcal{P}}_{\mathrm{GSS}}^{-1}\mathcal{N}_{\mathrm{GSS}}$ $=(\Theta+\bm{\omega}\mathcal{B})^{-1}(\Theta-(1-\bm{\omega})\mathcal{B})$ (4.1)
$=(I{+}\,\bm{\omega}\Theta^{-1}\mathcal{B})^{-1}(I-(1-\bm{\omega})\Theta^{-1}\mathcal{B}).$ (4.2)
@@
$(I{+}\,\bm{\omega}\Theta^{-1}\mathcal{B})^{-1}\left(I-(1-\bm{\omega})\Theta^{-1}\mathcal{B}\right){\bf x}=\lambda{\bf x},$ (4.3)
@@
$(I-(1-\bm{\omega})\Theta^{-1}\mathcal{B}){\bf x}=\lambda(I{+}\,\bm{\omega}\Theta^{-1}\mathcal{B}){\bf x}$ (4.4)
@@
Notably, solving the DSPP (1.1) is same as to solve the preconditioned linear system $\mathcal{P}_{\text{GSS}}^{-1}\mathcal{B}{\bf u}={\mathcal{P}}_{\text{GSS}}^{-1}{\bm{b}}$. Hence, as an immediate consequence of Theorem 4.1 and Remark 4.1, we have the following results regarding the clustering properties of the spectrum of the preconditioned matrix ${\mathcal{P}}_{\text{GSS}}^{-1}\mathcal{B}$.
@@
Assume that $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{p\times p}$ are nonsingular matrices with $A_{H}$ and $D_{H}$ positive definite, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are full row matrices. Let $\mathcal{P}_{\mathrm{GSS}}$ be defined as in (3.2) and $\lambda$ be an eigenvalue of the preconditioned matrix $\mathcal{P}_{\mathrm{GSS}}^{-1}\mathcal{B}$. If $\bm{\omega}\geq 1/2,$ then $\lambda$ satisfies the following:
@@
i.e., all eigenvalues of the preconditioned matrix $\mathcal{P}_{\mathrm{GSS}}^{-1}\mathcal{B}$ are entirely contained in a circle centered at $(1,0)$ with radius strictly less one.
@@
The proof follows immediately from the identity $\mathcal{P}_{\text{GSS}}^{-1}\mathcal{B}=\mathcal{P}_{\text{GSS}}^{-1}(\mathcal{P}_{\text{GSS}}-\mathcal{N}_{\text{GSS}})=I-\mathcal{G}.$blacksquare
@@ 5 Two relaxed variants of GSS preconditioner
To enhance efficiency, this section introduces two relaxed variants of the GSS preconditioner. By removing $\alpha P$ from $(1,1)$ block and $\beta Q$ from $(2,2)$ block of $\mathcal{P}_{\mathrm{GSS}},$ we obtain the following two relaxed GSS (RGSS) preconditioners:
$\mathcal{P}_{\text{RGSS-I}}=\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{\omega}B^{T}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{\omega}C\\ -\bm{\omega}B&-\bm{\omega}C^{T}&\tau R\end{matrix}\right]$ (5.1)
@@
$\mathcal{P}_{\text{RGSS-II}}=\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{\omega}B^{T}\\ \bm{0}&\bm{\omega}D&\bm{\omega}C\\ -\bm{\omega}B&-\bm{\omega}C^{T}&\tau R\end{matrix}\right].$ (5.2)
In the implementation, at each step of RGSS-I and RGSS-II preconditioners in conjunction with GMRES, we are required to solve the linear systems of the following forms:
$\mathcal{P}_{\text{RGSS-I}}{\bm{w}}={\bm{r}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{P}_{\text{RGSS-II}}{\bm{w}}={\bm{r}},$ (5.3)
respectively.
@@
$\mathcal{P}_{\text{RGSS-I}}=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -BA^{-1}&-\bm{\omega}C^{T}(\beta Q+\bm{\omega}D)^{-1}&I\end{matrix}\right]\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\mathcal{R}_{1}\end{matrix}\right]\left[\begin{matrix}I&\bm{0}&A^{-1}B^{T}\\ \bm{0}&I&\bm{\omega}(\beta Q+\bm{\omega}D)^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$ (5.4)
$\mathcal{P}_{\text{RGSS-II}}=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -BA^{-1}&-C^{T}D^{-1}&I\end{matrix}\right]\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\mathcal{R}_{2}\end{matrix}\right]\left[\begin{matrix}I&\bm{0}&A^{-1}B^{T}\\ \bm{0}&I&D^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$ (5.5)
By applying a similar methodology as in Algorithm 1, we derive Algorithms 2 and 3 for implementing the RGSS-I and RGSS-II preconditioners, respectively.
Algorithm 2 Solving $\mathcal{P}_{\text{RGSS-I}}{\bm{w}}={\bm{r}}$
Input: The matrices $A\in{\mathbb{R}}^{n\times n},\,B\in{\mathbb{R}}^{m\times n},\,C\in{\mathbb{R}}^{l\times m},\,D\in{\mathbb{R}}^{l\times l},$ $\bm{r}\in{\mathbb{R}}^{n+l+m},$ positive parameters $\beta,\tau,\bm{\omega},$ SPD matrices $Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m}.$
@@
$1:$ Solve $At_{1}=\bm{r}_{1}/\bm{\omega}$ to find $t_{1}.$
$2:$ Solve $(\beta Q+\bm{\omega}D)t_{2}=\bm{r}_{2}$ to find $t_{2}.$
$3:$ Solve $\mathcal{R}_{1}\bm{w}_{3}=\bm{r}_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2}$ to obtain $\bm{w}_{3}.$
$4:$ Solve $A\bm{w}_{1}=\frac{1}{\bm{\omega}}(\bm{r}_{1}-B^{T}\bm{w}_{3})$ to find $\bm{w}_{1}.$
$5:$ Solve $(\beta Q+\bm{\omega}D)\bm{w}_{2}=\bm{r}_{2}-\bm{\omega}C\bm{w}_{3}$ to obtain $\bm{w}_{2}.$
Algorithm 3 Solving $\mathcal{P}_{\text{RGSS-II}}{\bm{w}}={\bm{r}}$
Input: The matrices $A\in{\mathbb{R}}^{n\times n},\,B\in{\mathbb{R}}^{m\times n},\,C\in{\mathbb{R}}^{l\times m},\,D\in{\mathbb{R}}^{l\times l},$ $\bm{r}\in{\mathbb{R}}^{n+l+m},$ positive parameters $\tau,\bm{\omega}$ and the SPD matrix $R\in{\mathbb{R}}^{m\times m}.$
@@
$1:$ Solve $At_{1}=\bm{r}_{1}/\bm{\omega}$ to find $t_{1}.$
$2:$ Solve $Dt_{2}=\bm{r}_{2}/\bm{\omega}$ to find $t_{2}.$
$3:$ Solve $\mathcal{R}_{2}\bm{w}_{3}=\bm{r}_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2}$ to obtain $\bm{w}_{3}.$
$4:$ Solve $A\bm{w}_{1}=\frac{1}{\bm{\omega}}(\bm{r}_{1}-B^{T}\bm{w}_{3})$ to find $\bm{w}_{1}.$
$5:$ Solve $D\bm{w}_{2}=\frac{1}{\bm{\omega}}(\bm{r}_{2}-\bm{\omega}C\bm{w}_{3})$ to obtain $\bm{w}_{2}.$
Remark 5.1.
A key challenge in implementing the RGSS-I and RGSS-II preconditioners lies in solving the linear subsystems associated with the coefficient matrices $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$, respectively. As noted in Remark 3.2, to avoid the direct computation $BA^{-1}B^{T}$ and $C^{T}(\beta Q+\bm{\omega}D)^{-1}C$ in Algorithm 2, and $BA^{-1}B^{T}$ and $C^{T}D^{-1}C$ in Algorithm 3, we can use approximate versions of these terms. Following a similar technique as in decomposition (3.2), let $\widetilde{P}_{1},$ $\widetilde{Q}_{1}$ and $\widetilde{Q}_{2}$ be efficient and economical approximations of the matrices $BA^{-1}B^{T},$ $C^{T}(\beta Q+\bm{\omega}D)^{-1}C,$ and $C^{T}D^{-1}C,$ respectively. With these approximations, step $4$ in Algorithms 2 and 3 transforms into solving the linear subsystems $(\widetilde{P}_{1}+\widetilde{Q}_{1})\bm{z}_{3}={\bm{r}}_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2}$ and $(\widetilde{P}_{1}+\widetilde{Q}_{2})\bm{z}_{3}={\bm{r}}_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2},$ respectively. By appropriately selecting $\widetilde{P}_{1}$ $\widetilde{Q}_{1},$ and $\widetilde{Q}_{2}$ (see for Example $1$), these subsystems become significantly easier to implement compared to step $3$ of Algorithms 2 and 3. The resulting inexact preconditioners are denoted by $\widetilde{\mathcal{P}}_{\text{RGSS-I}}$ and $\widetilde{\mathcal{P}}_{\text{RGSS-II}},$ respectively.
Next, we investigate the spectral properties of the preconditioned matrices $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B},$ by considering $A$ and $D$ are SPD.